OpenStudy (anonymous):

Evaluate: 32 + (8 − 2) • 4 −6/3 30 58 28 31

jhonyy9 (jhonyy9):

so 8-2=6 and 6/3 =2 than how will be your exercise ? 32 +6*4 -2 =32 +24 -2 = 32 +22 = 54